# Measurement-based quantum computation utilizing the graph states of Bose-Einstein condensates and continuous variables

### v1

@@ Abstract
Measurement-based quantum computation (MBQC) is a protocol for quantum computation that represents a model distinct from the circuit-based approach. MBQC has been proposed not only for qubits but also for qudits, continuous-variable (CV qubits), and Bose-Einstein condensates (BECs) qubits. In qubit-based MBQC, arbitrary rotations on the Bloch sphere can be performed by measuring a graph state. This naturally raises the question of whether arbitrary rotations on the Bloch sphere can similarly be achieved through measurements in other types of quantum bits. We have demonstrated that this can indeed be realized for BECs qubits by considering composite graph states involving CV qubits and BECs qubits.
@@
In quantum computation protocols, Bose-Einstein condensates (BECs) quantum computing has been proposed 10 . Typically, when dealing with high-dimensional Hilbert spaces, such as those for qudits or qutrits, the Bloch sphere becomes more complex, making it visually challenging to ascertain whether arbitrary quantum computations are feasible. In the case of BECs qubits, it looks like, despite the high-dimensional Hilbert spaces, quantum states can be represented on a single Bloch sphere. This makes the consideration of arbitrary rotations on the Bloch sphere particularly meaningful. Also, BECs quantum computation has seen theoretical advancements in recent years 11 ; 12 ; 13 ; 14 ; 15 ; 16 , and particularly relevant to our study is the quantum repeater protocol 11 . In Ref 11 , $S^{z}S^{z}$ interactions are performed at a magic time to entangle atomic ensembles. Quantum repeaters are then achieved by measuring intermediate nodes. However, in the quantum repeater protocol, quantum operations achieved through measurements are constrained to a discrete range. This is because the measurement results yield a Kronecker delta function.
@@
In this paper, we theoretically realized arbitrary rotation on the BECs Bloch sphere using composite graph states consisting of BECs qubits and CV qubits. However, to achieve universal quantum computation, not only arbitrary rotations on the BECs Bloch sphere but also two-qubit gates, such as the CZ gate, are required. This will be discussed later. The fundamental idea is the same as in MBQC: prepare a graph state and perform quantum computation through measurements. What distinguishes this approach from other MBQC protocols is the consideration of graph states that mix two types of qubits, namely BECs qubits and CV qubits. As a result, continuous variables emerge, and it is noteworthy that Dirac delta functions arising due to measurements play a critical role. In the quantum repeater protocol 11 mentioned above, when measuring a intermediate node, Kronecker delta functions appear, leading to discrete quantum operations. In contrast, our framework enables continuous quantum operations by integrating over the CV qubits to resolve Dirac delta divergences. We also discussed the feasibility of arbitrary rotation on the Bloch sphere for representative graph structures.
@@ II BECs qubits and CV qubits
$|\alpha,\beta\rangle\rangle\equiv\frac{1}{\sqrt{N!}}(\alpha a^{\dagger}+\beta b^{\dagger})^{N}|\rm{vac}\rangle.$ (1)
@@
$\begin{split}S^{x}&=a^{\dagger}b+b^{\dagger}a\\ S^{y}&=-ia^{\dagger}b+ib^{\dagger}a\\ S^{z}&=a^{\dagger}a-b^{\dagger}b\\ N&=a^{\dagger}a+b^{\dagger}b\\ &=n^{a}+n^{b}.\end{split}$ (2)
@@ III.2 Measurement graph states and arbitrary rotation on the BECs Bloch sphere
where we choice the “magic time” time evolution $t=2\pi/L$ 11 between the CV qubit 2 and the BECs qubit 3, where we select the $L$ is integer and the large number and time evolution $t=1$ between BECs qubit 1 and CVqubit 2. On the other hand, the graph state, in the case of Fig. 3 (b), is
@@
$\begin{split}|\theta_{3}^{(x)}\rangle&=e^{-in^{b}_{3,l}(2\pi\theta/L-\pi/2)}e^{-in^{b}_{3,m}(2\pi\theta/L-\pi/2)}\\ &\times e^{-3iS^{y}\pi/4}|k\rangle.\end{split}$ (12)
We note that the state in which Hadamard gate $(H=e^{-3iS^{y}\pi/4})$ acts on the Fock basis is $(a^{\dagger}+b^{\dagger})^{k}(a^{\dagger}-b^{\dagger})^{N-k}/\sqrt{k!(N-k)!}|{\rm vac}\rangle$. $n^{b}_{3,l}$ and $n^{b}_{3,m}$ act as $n^{b}_{3,l}|l\rangle=(k-l)|l\rangle$, $n^{b}_{3,m}|m\rangle=(N-k-m)|m\rangle$, where $|l\rangle=(a^{\dagger})^{l}(b^{\dagger})^{k-l}/\sqrt{l!(k-l)!}|{\rm vac}\rangle$ and $|m\rangle=(a^{\dagger})^{m}(b^{\dagger})^{N-k-m}/\sqrt{m!(N-k-m)!}|{\rm vac}\rangle$. The results of the measurements for BECs 3 at this basis are
$\begin{split}&\langle\theta^{(x)}_{3}|\frac{1}{\sqrt{2}},\frac{e^{-2\pi ix/L}}{\sqrt{2}}\rangle\rangle_{3}\\ &=N!\cos^{q_{3}}(\pi(x-\theta)/L+\pi/4)\sin^{N-q_{3}}(\pi(x-\theta)/L+\pi/4)\\ &=N!\cos^{N}(\pi(x-\theta)/L+\pi/4)\tan^{N-q_{3}}(\pi(x-\theta)/L+\pi/4),\end{split}$ (13)
@@
which is valid for $\theta<<1$. In the eq. (14) case, $\pi(x-\theta)<<L$ is required. Therefore, the squared term of the exponential function in eq. (14) can be replaced with the Dirac delta function.
@@
Eq. (16) is valid for $\pi^{2}/2L^{2}<<N$. so, $N\rightarrow\infty$ is required. Therefore, the state after measuring $|G\rangle\rangle_{z}$ is
$\begin{split}|G\rangle\rangle_{z}^{\prime}&=N!\sqrt{\frac{L}{N}}\int_{-\infty}^{\infty}|\frac{1}{\sqrt{2}},\frac{e^{-ix}}{\sqrt{2}}\rangle\rangle_{1}\psi(x)|x\rangle_{2}\delta(x-\theta)\\ &\times\tan^{N-q_{3}}(\pi(x-\theta)/L+\pi/4)\exp(\pi(x-\theta)/L)dx.\end{split}$ (18)
@@
$|G\rangle\rangle_{z}^{\prime}=N!\sqrt{\frac{L}{N}}|\frac{1}{\sqrt{2}},\frac{e^{-i\theta}}{\sqrt{2}}\rangle\rangle_{1}\psi(\theta)|\theta\rangle_{2}.$ (19)

### v2

@@ Abstract
Measurement-based quantum computation (MBQC) is a protocol for quantum computation that represents a model distinct from the circuit-based approach. MBQC has been proposed not only for qubits but also for qudits, continuous-variable (CV) qubits, and Bose-Einstein condensates (BECs) qubits. In qubit-based MBQC, arbitrary rotations on the Bloch sphere can be performed by measuring a graph state. This naturally raises the question of whether arbitrary rotations on the Bloch sphere can similarly be achieved through measurements in other types of quantum bits. We have demonstrated that this can indeed be realized for BECs qubits by considering composite graph states involving CV qubits and BECs qubits.
@@
In quantum computation protocols, Bose-Einstein condensates (BECs) quantum computing has been proposed 10 . Typically, when dealing with high-dimensional Hilbert spaces, such as those for qudits or qutrits, the Bloch sphere becomes more complex, making it visually challenging to ascertain whether arbitrary Bloch sphere rotations are feasible. In the case of BECs qubits, it looks like, despite the high-dimensional Hilbert spaces, quantum states can be represented on a single Bloch sphere. This makes the consideration of arbitrary rotations on the Bloch sphere particularly meaningful. Also, BECs quantum computation has seen theoretical advancements in recent years 11 ; 12 ; 13 ; 14 ; 15 ; 16 , and particularly relevant to our study is the quantum repeater protocol 11 . In Ref 11 , $S^{z}S^{z}$ interactions are performed at a magic time to entangle atomic ensembles. The quantum repeater is then achieved by measuring intermediate nodes. However, in the quantum repeater protocol, quantum operations achieved through measurements are constrained to a discrete range. This is because the measurement results yield a Kronecker delta function.
@@
In this paper, we theoretically realized arbitrary rotation on the BECs Bloch sphere using composite graph states consisting of BECs qubits and CV qubits. However, to achieve universal quantum computation, not only arbitrary rotations on the BECs Bloch sphere but also two-qubit gates, such as the CZ gate, are required. This will be discussed later. The fundamental idea is the same as in MBQC: prepare a graph state and perform quantum computation through measurements. What distinguishes this approach from other MBQC protocols is the consideration of graph states that mix two types of qubits, namely BECs qubits and CV qubits. As a result, continuous variables emerge, and it is noteworthy that Dirac delta functions arising due to measurements play a critical role. In the quantum repeater protocol 11 mentioned above, when measuring a intermediate node, Kronecker delta functions appear, leading to discrete quantum operations. In contrast, our framework enables continuous quantum operations by integrating over the CV qubits to resolve Dirac delta divergences. We will also discuss the feasibility of arbitrary rotation on the Bloch sphere for representative graph structures.
@@ II BECs qubits and CV qubits
$|\alpha,\beta\rangle\rangle\equiv\frac{1}{\sqrt{N!}}(\alpha a^{\dagger}+\beta b^{\dagger})^{N}|\rm{vac}\rangle,$ (1)
@@
$\begin{split}S^{x}&=a^{\dagger}b+b^{\dagger}a,\\ S^{y}&=-ia^{\dagger}b+ib^{\dagger}a,\\ S^{z}&=a^{\dagger}a-b^{\dagger}b,\\ N&=a^{\dagger}a+b^{\dagger}b\\ &=n^{a}+n^{b}.\end{split}$ (2)
@@ III.2 Measurement graph states and arbitrary rotation on the BECs Bloch sphere
where we choice the “magic time” time evolution $t=2\pi/L$ 11 between the CV qubit 2 and the BECs qubit 3, where we select the $L$ is integer and the large number and time evolution $t=1$ between BECs qubit 1 and CVqubit 2. On the other hand, the graph state, in the case of Fig. 3 (b), is [Uncaptioned image]
@@
$\begin{split}|\theta_{3}^{(x)}\rangle&=e^{-in^{b}_{3,l}(2\pi\theta/L-\pi/2)}e^{-in^{b}_{3,m}(2\pi\theta/L-\pi/2)}\\ &\times e^{-3iS^{y}\pi/4}|k\rangle,\end{split}$ (12)
which is an unnormalized state. We note that the state in which Hadamard gate $(H=e^{-3iS^{y}\pi/4})$ acts on the Fock basis is $(a^{\dagger}+b^{\dagger})^{k}(a^{\dagger}-b^{\dagger})^{N-k}/\sqrt{k!(N-k)!}|{\rm vac}\rangle$. $n^{b}_{3,l}$ and $n^{b}_{3,m}$ act as $n^{b}_{3,l}|l\rangle=(k-l)|l\rangle$, $n^{b}_{3,m}|m\rangle=(N-k-m)|m\rangle$, where $|l\rangle=(a^{\dagger})^{l}(b^{\dagger})^{k-l}/\sqrt{l!(k-l)!}|{\rm vac}\rangle$ and $|m\rangle=(a^{\dagger})^{m}(b^{\dagger})^{N-k-m}/\sqrt{m!(N-k-m)!}|{\rm vac}\rangle$. The results of the measurements for BECs 3 at this basis are
$\begin{split}&\langle\theta^{(x)}_{3}|\frac{1}{\sqrt{2}},\frac{e^{-2\pi ix/L}}{\sqrt{2}}\rangle\rangle_{3}\\ &=\sqrt{\frac{N!}{q_{3}!(N-q_{3})!}}\cos^{q_{3}}(\pi(x-\theta)/L+\pi/4)\\ &\times\sin^{N-q_{3}}(\pi(x-\theta)/L+\pi/4)\\ &=\sqrt{\frac{N!}{q_{3}!(N-q_{3})!}}\cos^{N}(\pi(x-\theta)/L+\pi/4)\\ &\times\tan^{N-q_{3}}(\pi(x-\theta)/L+\pi/4),\end{split}$ (13)
@@
which is valid for $\theta\ll 1$. In the eq. (14) case, $\pi(x-\theta)\ll L$ is required. Therefore, the squared term of the exponential function in eq. (14) can be replaced with the Dirac delta function.
@@
Eq. (16) is valid for $\pi^{2}/2L^{2}\ll N$. so, $N\rightarrow\infty$ is required. Therefore, the state after measuring $|G\rangle\rangle_{z}$ is
$\begin{split}|G\rangle\rangle_{z}^{\prime}&=\sqrt{\frac{N!}{q_{3}!(N-q_{3})!}}\sqrt{\frac{L}{2^{N}N}}\int_{-\infty}^{\infty}|\frac{1}{\sqrt{2}},\frac{e^{-ix}}{\sqrt{2}}\rangle\rangle_{1}\\ &\times\psi(x)|x\rangle_{2}\delta(x-\theta)\tan^{N-q_{3}}(\pi(x-\theta)/L+\pi/4)\\ &\times\exp(\pi(x-\theta)/L)dx.\end{split}$ (18)
@@
$|G\rangle\rangle_{z}^{\prime}=\sqrt{\frac{N!}{q_{3}!(N-q_{3})!}}\sqrt{\frac{L}{2^{N}N}}|\frac{1}{\sqrt{2}},\frac{e^{-i\theta}}{\sqrt{2}}\rangle\rangle_{1}\psi(\theta)|\theta\rangle_{2}.$ (19)